# Community-Based Early-Stage Chronic Kidney Disease Screening using Explainable Machine Learning for Low-Resource Settings

### v1

@@ 2.2 Data Preprocessing
Among the four datasets incorporated into this study, the UCI-2015 CKD dataset [chronic_kidney_disease_336] contained missing values requiring systematic preprocessing. Missing numeric features were addressed using the multivariate imputation by chained equations (MICE) algorithm, a widely adopted multiple-imputation framework that iteratively models each variable with missing data as a function of the other variables in the dataset [vanBuuren2011MICE]. Rather than replacing missing entries with a single deterministic value, MICE generates multiple plausible imputations, thereby preserving the underlying variability and reducing bias associated with single-imputation techniques. For categorical variables, missing entries were imputed using the most-frequent strategy, an approach demonstrated to be effective for maintaining distributional integrity in classification-focused machine learning tasks [adhikari2022comprehensive].
To ensure consistency, shared features across external and primary datasets were identified, harmonized, and renamed. Continuous variables were discretized into categorical, where appropriate, following our dataset’s strategy (Table 1). This normalization aligned heterogeneous sources, enabling model training on the primary dataset and evaluation on external datasets.
Finally, all four datasets were transformed using one-hot encoding to convert nominal categorical variables into a numerical format compatible with machine learning algorithms [pedregosa2011scikit]. This approach avoids spurious ordinal assumptions, improves interpretability, and ensures consistency across the diverse classification models evaluated in this study.
@@ 2.3 Feature Ranking and Selection
Beyond group-wise evaluation, we employed a diverse set of feature selection (FS) techniques to identify the most informative subset of predictors across all features (denoted as set S1) and across all features except the pathology tests group (denoted as set S2). Ten complementary FS methods were applied, including a statistical filter technique such as Mann–Whitney U (MWU) test [mann1947test], a regularization-based method such as the least absolute shrinkage and selection operator (LASSO) [tibshirani1996lasso], and wrapper-based recursive approaches such as recursive feature elimination with cross-validation (RFECV) [guyon2002rfe] using a variety of ML estimators, including logistic regression (LR) [hosmer2013applied], decision tree (DT) [breiman1984cart], random forest (RF) [breiman2001rf], gradient boosting (GB) [friedman2001gbm], adaptive boosting (AB) [freund1997adaboost], extra trees (ET) [geurts2006extratrees], extreme gradient boosting (XGB) [chen2016xgboost], and CatBoost (CB) [dorogush2018catboost].
@@ 2.4 Machine Learning Classifiers
To develop a robust and generalizable model for early CKD prediction, we employed twelve state-of-the-art ML classifiers spanning diverse algorithmic paradigms. These classifiers were selected based on their demonstrated effectiveness in handling structured, tabular, and high-dimensional clinical data, as well as their widespread use in prior CKD detection research [sanmarchi2023predict, delrue2024application, gogoi2025machine]. The models encompass linear, nonlinear, ensemble-based, and neural network–based learning strategies, thereby ensuring comprehensive exploration of the feature space.
@@
The set includes a linear model such as logistic regression (LR) [hosmer2013applied], a tree-based classifier such as decision tree (DT) [breiman1984cart], a distance-based model such as $k$-nearest neighbours (KNN) [cover1967knn], and a kernel-based algorithm such as support vector machine (SVM) [cortes1995svm]. To capture variance via bootstrap aggregation (bagging), we incorporated random forest (RF) [breiman2001rf] and extra trees (ET) [geurts2006extratrees]. For sequential error-correcting boosting strategies, we employed adaptive boosting (AB) [freund1997adaboost], gradient boosting (GB) [friedman2001gbm], extreme gradient boosting (XGB) [chen2016xgboost], light gradient boosting (LGB) [ke2017lightgbm], and CatBoost (CB) [dorogush2018catboost]. Finally, we included a neural network–based classifier, the multi-layer perceptron (MLP) [rumelhart1986mlp], to evaluate nonlinear representations learned through back-propagation. Together, these twelve classifiers provide a comprehensive benchmark for assessing CKD predictability across various modeling assumptions and complexity levels.
2.5 Model Training
@@
All machine learning models were trained following a rigorous and reproducible experimental pipeline. To ensure robust estimation of generalization performance, we employed stratified 10-fold cross-validation, a widely used model evaluation strategy that mitigates overfitting by repeatedly training and testing models on different partitions of the data while preserving class distribution [kohavi1995cv]. For reproducibility, all random processes (data shuffling, model initialization, and parameter sampling) were controlled using a fixed random_state = 42.
@@
Model development and evaluation were implemented using the scikit-learn library [pedregosa2011scikit], along with specialized optimization frameworks for advanced gradient boosting models. To identify the optimal hyperparameters for each classifier, we utilized Optuna, an efficient, modern hyperparameter optimization framework that employs sequential model-based optimization (SMBO) with adaptive sampling strategies to search high-dimensional parameter spaces [akiba2019optuna]. During optimization, each candidate configuration was evaluated within the cross-validation workflow to ensure unbiased performance estimation. This systematic training and optimization procedure enabled the selection of high-performing, well-calibrated models across all feature subsets and ensured fair comparison among the twelve machine learning classifiers.
2.6 Performance Metrics
@@
Given that the dataset used in this study is imbalanced, with the CKD class representing the minority and clinically significant positive class, we employed evaluation metrics that are robust to class imbalance and that reflect clinically meaningful performance. For binary classification (CKD as the positive class), we report balanced accuracy, sensitivity (recall for CKD), area under the receiver operating characteristic curve (AUC-ROC), F1 score (for CKD), and macro-precision. The definition of those metrics is provided in B. For all metrics, we additionally compute the 95% confidence intervals (CI) across the 10-fold cross-validation splits to quantify statistical uncertainty. This provides a robust estimate of model reliability and generalization performance.
@@ 2.7 Explainability
Explainability is a critical component in the development and deployment of machine learning models for clinical applications. It enhances transparency, fosters trust among healthcare professionals, and supports clinical validation by elucidating how input features contribute to model predictions. In the context of CKD detection, explainability enables clinicians to understand which factors most strongly influence an individual’s predicted CKD risk, thereby facilitating more informed decision-making and improving patient management.
Among the different explainability methods, SHAP (SHapley Additive exPlanations) [lundberg2017shap] has emerged as one of the most powerful due to its rigorous theoretical grounding in cooperative game theory. SHAP provides both global interpretability by quantifying overall feature importance across the population and local interpretability, allowing an examination of how specific features contribute to the prediction for an individual patient. These dual capabilities make SHAP especially suitable for clinical environments, where understanding both broad model behaviour and case-specific reasoning is essential. SHAP has also been successfully applied in previous studies on CKD prediction [gogoi2025machine].
@@ 3.1 Selected Features
3.2 Performance of ML Models
To assess the discriminative capability of different feature domains for CKD prediction, we systematically evaluated multiple feature sets, including individual feature groups, their combinations, and the feature subsets derived from S1 and S2 using the ten feature selection techniques described in Section 3.1. This comprehensive comparison enables a detailed understanding of how various domains and selected features contribute to model performance.
Table 3 summarizes the performance of all feature sets, reporting results based on the best-performing model for each set. Every feature set was evaluated independently using the twelve machine learning classifiers introduced in Section 2.4. All models were trained and validated using stratified 10-fold cross-validation, and performance was quantified using appropriate evaluation metrics accompanied by 95% confidence intervals to ensure statistical robustness.
The best-performing model for each feature set was identified primarily based on the highest balanced accuracy, reflecting the need to account for class imbalance in CKD prediction. In cases where two or more models achieved identical balanced accuracy, we applied a secondary ranking strategy prioritizing higher sensitivity for the CKD (positive) class, followed by AUC-ROC, and subsequently the CKD-specific F1 score. This hierarchical selection process ensures that the chosen model not only performs well overall but also aligns with clinical priorities by maximizing the correct identification of CKD cases.
@@ 3.2 Performance of ML Models
Table 3: Performance of best models for various feature sets. Bold values represent the highest performance metric in each group
Feature set Best Model Performance $\pm$ 95% Confidence Interval (CI)
@@
All CB $84.54_{\pm 3.27}$ $79.55_{\pm 5.94}$ $0.900_{\pm 0.05}$ $81.22_{\pm 4.00}$ $85.30_{\pm 2.91}$
Groups and combinations Socio-Demo (SD) DT $73.27_{\pm 2.75}$ $65.23_{\pm 6.90}$ $0.816_{\pm 0.03}$ $67.08_{\pm 3.69}$ $74.35_{\pm 2.88}$
Life-style/Habbit (LH) LGB $62.90_{\pm 6.23}$ $53.64_{\pm 10.19}$ $0.648_{\pm 0.07}$ $54.09_{\pm 8.20}$ $63.18_{\pm 6.00}$
Medical-history (MH) DT $78.87_{\pm 7.09}$ $75.15_{\pm 9.74}$ $0.805_{\pm 0.06}$ $74.37_{\pm 8.40}$ $78.99_{\pm 7.13}$
Clinical exam (CE) LR $58.77_{\pm 6.48}$ $57.12_{\pm 12.50}$ $0.593_{\pm 0.06}$ $51.68_{\pm 8.62}$ $58.59_{\pm 6.37}$
Pathology tests (Path) LR $64.24_{\pm 5.74}$ $52.80_{\pm 10.14}$ $0.702_{\pm 0.07}$ $55.12_{\pm 7.41}$ $66.13_{\pm 6.51}$
SD-LH AB $73.20_{\pm 2.98}$ $49.92_{\pm 6.45}$ $0.820_{\pm 0.04}$ $63.95_{\pm 5.31}$ $\textbf{83.53}_{\pm 3.95}$
SD-LH-MH LR $\textbf{82.93}_{\pm 3.55}$ $\textbf{78.64}_{\pm 6.71}$ $0.881_{\pm 0.04}$ $\textbf{79.16}_{\pm 4.26}$ $83.40_{\pm 3.22}$
SD-LH-MH-CE CB $81.87_{\pm 4.26}$ $74.17_{\pm 6.72}$ $\textbf{0.887}_{\pm 0.06}$ $77.80_{\pm 5.59}$ $83.21_{\pm 3.96}$
Selected across all features (S1) RFECV+LR AB $86.19_{\pm 3.67}$ $84.02_{\pm 8.40}$ $0.89_{\pm 0.04}$ $82.91_{\pm 4.30}$ $86.35_{\pm 2.74}$
RFECV+RF SVM $88.45_{\pm 5.17}$ $76.89_{\pm 10.34}$ $0.859_{\pm 0.08}$ $86.24_{\pm 6.80}$ $\textbf{93.72}_{\pm 2.56}$
RFECV+GB GB $88.75_{\pm 5.50}$ $78.64_{\pm 10.31}$ $0.920_{\pm 0.07}$ $86.56_{\pm 7.17}$ $92.96_{\pm 3.70}$
RFECV+DT AB $86.19_{\pm 3.67}$ $84.02_{\pm 8.40}$ $0.895_{\pm 0.04}$ $82.91_{\pm 4.30}$ $86.35_{\pm 2.74}$
RFECV+AB AB $86.19_{\pm 3.67}$ $84.02_{\pm 8.40}$ $0.895_{\pm 0.04}$ $82.91_{\pm 4.30}$ $86.35_{\pm 2.74}$
RFECV+ET GB $87.99_{\pm 4.67}$ $75.98_{\pm 9.34}$ $\textbf{0.931}_{\pm 0.04}$ $85.76_{\pm 6.32}$ $93.44_{\pm 2.23}$
RFECV+XGB XGB $88.16_{\pm 6.17}$ $80.38_{\pm 10.77}$ $0.915_{\pm 0.07}$ $85.76_{\pm 7.85}$ $91.11_{\pm 5.32}$
RFECV+CB DT $\textbf{90.40}_{\pm 5.07}$ $\textbf{84.85}_{\pm 7.39}$ $0.906_{\pm 0.05}$ $\textbf{88.77}_{\pm 6.43}$ $92.14_{\pm 4.99}$
LASSO LR $84.99_{\pm 3.82}$ $82.20_{\pm 9.09}$ $0.910_{\pm 0.04}$ $81.46_{\pm 4.40}$ $85.56_{\pm 3.17}$
MWUtest GB $85.99_{\pm 4.26}$ $79.47_{\pm 7.49}$ $0.925_{\pm 0.05}$ $83.09_{\pm 5.42}$ $87.91_{\pm 4.38}$
Union CB $85.88_{\pm 3.47}$ $80.45_{\pm 6.47}$ $0.898_{\pm 0.05}$ $82.88_{\pm 4.35}$ $87.00_{\pm 3.24}$
Common AB $86.19_{\pm 3.67}$ $84.02_{\pm 8.40}$ $0.895_{\pm 0.04}$ $82.91_{\pm 4.30}$ $86.35_{\pm 2.74}$
Selected across all except pathology tests (S2) RFECV+LR AB $83.50_{\pm 3.53}$ $78.64_{\pm 7.38}$ $0.861_{\pm 0.04}$ $79.77_{\pm 4.41}$ $84.21_{\pm 2.80}$
RFECV+RF MLP $86.25_{\pm 4.79}$ $77.73_{\pm 10.57}$ $0.914_{\pm 0.05}$ $83.14_{\pm 5.92}$ $89.56_{\pm 3.50}$
RFECV+GB CB $\textbf{89.23}_{\pm 6.08}$ $\textbf{83.11}_{\pm 9.13}$ $0.906_{\pm 0.07}$ $\textbf{87.12}_{\pm 7.60}$ $\textbf{90.92}_{\pm 5.74}$
RFECV+DT DT $87.75_{\pm 7.44}$ $79.55_{\pm 11.52}$ $0.913_{\pm 0.07}$ $85.11_{\pm 9.57}$ $90.19_{\pm 6.77}$
RFECV+AB AB $83.50_{\pm 3.53}$ $78.64_{\pm 7.38}$ $0.861_{\pm 0.04}$ $79.77_{\pm 4.41}$ $84.21_{\pm 2.80}$
RFECV+ET CB $86.25_{\pm 4.79}$ $77.73_{\pm 10.57}$ $0.909_{\pm 0.05}$ $83.14_{\pm 5.92}$ $89.56_{\pm 3.50}$
RFECV+XGB RF $87.17_{\pm 7.05}$ $79.55_{\pm 11.52}$ $\textbf{0.921}_{\pm 0.06}$ $84.24_{\pm 8.99}$ $89.30_{\pm 6.27}$
RFECV+CB MLP $86.25_{\pm 4.79}$ $77.73_{\pm 10.57}$ $0.914_{\pm 0.05}$ $83.14_{\pm 5.92}$ $89.56_{\pm 3.50}$
LASSO LR $83.50_{\pm 3.53}$ $78.64_{\pm 7.38}$ $0.892_{\pm 0.04}$ $79.77_{\pm 4.41}$ $84.21_{\pm 2.80}$
MWUtest CB $84.42_{\pm 6.61}$ $78.64_{\pm 10.07}$ $0.896_{\pm 0.07}$ $80.84_{\pm 7.74}$ $85.61_{\pm 5.63}$
Union LGB $84.58_{\pm 6.60}$ $79.62_{\pm 11.08}$ $0.898_{\pm 0.06}$ $80.95_{\pm 8.19}$ $85.52_{\pm 5.95}$
Common AB $83.50_{\pm 3.53}$ $78.64_{\pm 7.38}$ $0.861_{\pm 0.04}$ $79.77_{\pm 4.41}$ $84.21_{\pm 2.80}$
@@
Overall, the results in Table 3 highlight that models trained on the selected features from the full feature set (S1) achieve the strongest predictive performance. The highest balanced accuracy of 90.40% is obtained using a DT model applied to the RFECV with CatBoost selected S1 feature subset (Table 1), achieving the highest sensitivity of 84.85% and F1 score of 88.77%. Among the individual feature groups, the Medical History (MH) and Socio-Demographic (SD) sets demonstrate competitive predictive ability, with balanced accuracies of 78.87% and 73.27%, respectively. Notably, the combined SD-LH-MH set outperforms all individuals and combined groups, achieving a balanced accuracy of 82.93%, which is higher than when clinical examination features are additionally included (i.e., SD-LH-MH-CE).

### v2

@@ 3.2 Data Preprocessing
Among the four datasets incorporated into this study, the UCI-2015 CKD dataset [64] contained missing values requiring systematic preprocessing. Missing numeric features were addressed using the multivariate imputation by chained equations (MICE) algorithm, a widely adopted multiple-imputation framework that iteratively models each variable with missing data as a function of the other variables in the dataset [73]. Rather than replacing missing entries with a single deterministic value, MICE generates multiple plausible imputations, thereby preserving the underlying variability and reducing bias associated with single-imputation techniques. For categorical variables, missing entries were imputed using the most-frequent strategy, an approach demonstrated to be effective for maintaining distributional integrity in classification-focused machine learning tasks [2].
To ensure consistency, shared features across external and primary datasets were identified, harmonized, and renamed. Continuous variables were discretized into categorical, where appropriate, following clinically grounded categories defined in the original cohort study (outlined in Table 2). Using these clinically grounded thresholds makes the resulting models more interpretable and better aligned with how risk factors are assessed in community health settings. Additionally, discretization helps reduce sensitivity to measurement noise and can improve model robustness when working with relatively small datasets. This discretization process aligned heterogeneous sources, enabling model training on the primary dataset and evaluation on external datasets.
Finally, all four datasets were transformed using one-hot encoding to convert nominal categorical variables into a numerical format compatible with machine learning algorithms [59]. This approach avoids spurious ordinal assumptions, improves interpretability, and ensures consistency across the diverse classification models evaluated in this study.
@@ 3.3 Feature Ranking and Selection
Beyond group-wise evaluation, we employed a diverse set of feature selection (FS) techniques to identify the most informative subset of predictors across all features (denoted as set S1) and across all features except the pathology tests group (denoted as set S2). Ten complementary FS methods were applied, including a statistical filter technique such as Mann–Whitney U (MWU) test [52], a regularization-based method such as the least absolute shrinkage and selection operator (LASSO) [72], and wrapper-based recursive approaches such as recursive feature elimination with cross-validation (RFECV) [25] using a variety of ML estimators, including logistic regression (LR) [27], decision tree (DT) [10], random forest (RF) [11], gradient boosting (GB) [20], adaptive boosting (AB) [19], extra trees (ET) [21], extreme gradient boosting (XGB) [12], and CatBoost (CB) [18].
@@
Feature ranking and selection were performed on the full primary dataset (TangailBD) prior to model training to identify stable predictor subsets suitable for community-based screening. Ten complementary feature selection techniques were applied to evaluate the importance of candidate variables and derive optimized feature subsets. Two subsets were constructed: S1, selected from all available variables, and S2, selected from variables excluding pathology-test features. These subsets were subsequently used for model development and evaluation. Because the goal of this step was to identify clinically meaningful predictors rather than tune model parameters, feature selection was performed once on the full dataset. The predictive performance of models built on these subsets was then assessed using nested cross-validation on the primary dataset and external validation on independent datasets.
@@ 3.4 Machine Learning Classifiers
To develop a robust and generalizable model for early CKD prediction, we employed twelve state-of-the-art ML classifiers spanning diverse algorithmic paradigms. These classifiers were selected based on their demonstrated effectiveness in handling structured, tabular, and high-dimensional clinical data, as well as their widespread use in prior CKD detection research [67, 16, 24]. The models encompass linear, nonlinear, ensemble-based, and neural network–based learning strategies, thereby ensuring comprehensive exploration of the feature space.
@@
The set includes a linear model such as logistic regression (LR) [27], a tree-based classifier such as decision tree (DT) [10], a distance-based model such as $k$-nearest neighbours (KNN) [14], and a kernel-based algorithm such as support vector machine (SVM) [13]. To capture variance via bootstrap aggregation (bagging), we incorporated random forest (RF) [11] and extra trees (ET) [21]. For sequential error-correcting boosting strategies, we employed adaptive boosting (AB) [19], gradient boosting (GB) [20], extreme gradient boosting (XGB) [12], light gradient boosting (LGB) [41], and CatBoost (CB) [18]. Finally, we included a neural network–based classifier, the multi-layer perceptron (MLP) [65], to evaluate nonlinear representations learned through back-propagation. Together, these twelve classifiers provide a comprehensive benchmark for assessing CKD predictability across various modeling assumptions and complexity levels.
3.5 Model Training and Evaluation
@@
All machine learning models were trained following a rigorous and reproducible experimental pipeline. To ensure robust estimation of generalization performance, we employed stratified 10-fold cross-validation, a widely used model evaluation strategy that mitigates overfitting by repeatedly training and testing models on different partitions of the data while preserving class distribution [45]. For reproducibility, all random processes (data shuffling, model initialization, and parameter sampling) were controlled using a fixed random_state = 42. To further assess model generalizability, external validation was subsequently performed using three independent datasets (UCI-2023, UCI-2015, and TH).
@@
Model development and evaluation were implemented using the scikit-learn library [59], along with specialized optimization frameworks for advanced gradient boosting models. To identify the optimal hyperparameters for each classifier, we utilized Optuna, an efficient, modern hyperparameter optimization framework that employs sequential model-based optimization (SMBO) with adaptive sampling strategies to search high-dimensional parameter spaces [3]. Hyperparameter optimization was conducted within a nested cross-validation framework. An outer stratified 10-fold cross-validation was used to estimate model performance, while hyperparameters were tuned within each training fold using an inner 5-fold cross-validation. This design helps prevent information leakage during hyperparameter tuning and provides an unbiased estimate of generalization performance. This systematic training and optimization procedure enabled the selection of high-performing, well-calibrated models across all feature subsets and ensured fair comparison among the twelve machine learning classifiers.
Model stability and variance were evaluated using 10-fold cross-validation. For each fold, models were trained on 90% of the data and evaluated on the held-out fold, producing out-of-fold predictions for all samples. Performance metrics were computed for each test fold and averaged across the 10 folds to obtain overall estimates of generalization performance. To quantify uncertainty and variability, 95% confidence intervals were estimated using stratified nonparametric bias-corrected and accelerated (BCa) bootstrap resampling (10,000 repetitions) of the pooled out-of-fold predictions. Tree-based and boosting models can exhibit variability due to their dependence on training data composition; however, cross-validation and bootstrap confidence intervals provide an empirical assessment of this variability and help ensure that reported performance reflects stable model behavior. To assess whether the performance difference between two models was statistically significant, we conducted a two-sided paired permutation test with 10,000 permutations using per-sample Brier loss calculated on the same evaluation samples for both models, with statistical significance assessed at $\alpha=0.05$.
@@
Given that the dataset used in this study is imbalanced, with the CKD class representing the minority and clinically significant positive class, we employed evaluation metrics that are robust to class imbalance and that reflect clinically meaningful performance. For binary classification (CKD as the positive class), we report balanced accuracy, sensitivity (recall for CKD), area under the receiver operating characteristic curve (AUC-ROC), F1 score (for CKD), and macro-precision. The definition of those metrics is provided in A. For binary classification, predicted probabilities were converted to CKD/non-CKD labels using a fixed decision threshold of 0.5, which was applied consistently across the primary dataset and all external validation datasets to ensure comparability of performance metrics.
The best-performing model for each feature set was selected primarily based on the highest balanced accuracy, reflecting the need to account for class imbalance in CKD prediction. In cases where two or more models achieved identical balanced accuracy, we applied a secondary ranking strategy prioritizing higher sensitivity for the CKD (positive) class, followed by AUC-ROC, and subsequently the CKD-specific F1 score. This hierarchical selection process ensures that the chosen model not only performs well overall but also aligns with clinical priorities by maximizing the correct identification of CKD cases.
@@ 3.6 Explainability
Among the different explainability methods, SHAP (SHapley Additive exPlanations) [50] has emerged as one of the most powerful due to its rigorous theoretical grounding in cooperative game theory. SHAP provides both global interpretability by quantifying overall feature importance across the population and local interpretability, allowing an examination of how specific features contribute to the prediction for an individual patient. These dual capabilities make SHAP especially suitable for clinical environments, where understanding both broad model behaviour and case-specific reasoning is essential. SHAP has also been successfully applied in previous studies on CKD prediction [24].
@@ 4.1 Selected Features
4.2 Performance of ML Models
Table 5 summarizes the performance of all feature sets, reporting results based on the best-performing model for each set. Every feature set was evaluated independently using the twelve machine learning classifiers introduced in Section 3.4.
Table 5: Predictive performance of the best machine learning models for different feature groups, feature combinations, and feature-selection. Results are reported as mean performance across 10-fold cross-validation with 95% confidence interval estimated using stratified nonparametric BCa bootstrap resampling of pooled fold predictions. The best-performing classifier for each feature set is reported. Bold values represent the highest performance metric in each group.
Feature set Best Model Performance with 95% confidence interval (CI)
@@ 4.2 Performance of ML Models
All CB $84.54_{\text{79.6--88.5}}$ $79.55_{\text{69.6--85.7}}$ $0.900_{\text{0.85--0.93}}$ $81.22_{\text{75.2--86.2}}$ $85.30_{\text{80.1--88.9}}$
Groups and combinations Socio-Demo (SD) DT $73.27_{\text{67.7--78.3}}$ $65.23_{\text{54.5--72.3}}$ $0.816_{\text{0.75--0.86}}$ $67.08_{\text{60.1--73.8}}$ $74.35_{\text{68.2--79.0}}$
Life-style/Habbit (LH) LGB $62.90_{\text{57.2--68.5}}$ $53.64_{\text{43.8--61.6}}$ $0.648_{\text{0.53--0.67}}$ $54.09_{\text{47.2--61.9}}$ $63.18_{\text{57.2--68.6}}$
Medical-history (MH) DT $78.87_{\text{73.5--83.2}}$ $75.15_{\text{65.2--81.3}}$ $0.805_{\text{0.73--0.85}}$ $74.37_{\text{67.9--79.8}}$ $78.99_{\text{73.1--83.0}}$
Clinical exam (CE) LR $58.77_{\text{52.9--64.6}}$ $57.12_{\text{47.3--65.2}}$ $0.593_{\text{0.51--0.65}}$ $51.68_{\text{45.6--59.0}}$ $58.59_{\text{52.8--64.0}}$
Pathology tests (Path) LR $64.24_{\text{58.3--69.6}}$ $52.80_{\text{42.9--60.7}}$ $0.702_{\text{0.60--0.73}}$ $55.12_{\text{47.5--62.8}}$ $66.13_{\text{58.7--70.5}}$
SD-LH AB $73.20_{\text{68.2--77.9}}$ $49.92_{\text{39.3--58.0}}$ $0.820_{\text{0.76--0.86}}$ $63.95_{\text{55.4--72.0}}$ $\textbf{83.53}_{\text{77.0--86.3}}$
SD-LH-MH LR $\textbf{82.93}_{\text{77.9--87.0}}$ $\textbf{78.64}_{\text{68.8--84.8}}$ $0.881_{\text{0.83--0.92}}$ $\textbf{79.16}_{\text{73.2--84.4}}$ $83.40_{\text{78.2--87.2}}$
SD-LH-MH-CE CB $81.87_{\text{76.9--86.2}}$ $74.17_{\text{64.3--81.3}}$ $\textbf{0.887}_{\text{0.84--0.93}}$ $77.80_{\text{71.5--83.5}}$ $83.21_{\text{78.2--87.3}}$
Selected across all features (S1) RFECV+LR AB $86.19_{\text{81.4--90.0}}$ $84.02_{\text{75.0--89.3}}$ $0.895_{\text{0.84--0.92}}$ $82.91_{\text{77.6--87.8}}$ $86.35_{\text{81.2--89.7}}$
RFECV+RF SVM $88.45_{\text{83.5--91.5}}$ $76.89_{\text{67.0--83.0}}$ $0.859_{\text{0.78--0.90}}$ $86.24_{\text{80.2--90.7}}$ $\textbf{93.72}_{\text{91.1--95.0}}$
RFECV+GB GB $88.75_{\text{84.4--92.1}}$ $78.64_{\text{68.8--84.8}}$ $0.920_{\text{0.85--0.94}}$ $86.56_{\text{81.4--91.3}}$ $92.96_{\text{89.5--94.8}}$
RFECV+DT AB $86.19_{\text{81.4--90.0}}$ $84.02_{\text{75.0--89.3}}$ $0.895_{\text{0.84--0.92}}$ $82.91_{\text{77.6--87.8}}$ $86.35_{\text{81.2--89.7}}$
RFECV+AB AB $86.19_{\text{81.4--90.0}}$ $84.02_{\text{75.0--89.3}}$ $0.895_{\text{0.84--0.92}}$ $82.91_{\text{77.6--87.8}}$ $86.35_{\text{81.2--89.7}}$
RFECV+ET GB $87.99_{\text{83.0--91.1}}$ $75.98_{\text{66.1--82.1}}$ $\textbf{0.931}_{\text{0.87--0.95}}$ $85.76_{\text{79.6--90.2}}$ $93.44_{\text{91.0--94.8}}$
RFECV+XGB XGB $88.16_{\text{83.5--91.8}}$ $80.38_{\text{70.5--85.7}}$ $0.915_{\text{0.86--0.95}}$ $85.76_{\text{80.2--90.7}}$ $91.11_{\text{86.3--93.4}}$
RFECV+CB DT $\textbf{90.40}_{\text{86.1--93.5}}$ $\textbf{84.85}_{\text{75.9--90.2}}$ $0.907_{\text{0.86--0.94}}$ $\textbf{88.77}_{\text{83.5--92.6}}$ $92.14_{\text{87.9--94.7}}$
LASSO LR $84.99_{\text{80.1--88.9}}$ $82.20_{\text{72.3--87.5}}$ $0.910_{\text{0.87--0.94}}$ $81.46_{\text{76.0--86.5}}$ $85.56_{\text{80.1--88.7}}$
MWUtest GB $85.99_{\text{81.2--89.8}}$ $79.47_{\text{69.6--85.7}}$ $0.925_{\text{0.87--0.95}}$ $83.09_{\text{77.3--88.0}}$ $87.91_{\text{82.6--90.8}}$
Union CB $85.88_{\text{81.1--89.6}}$ $80.45_{\text{71.4--85.7}}$ $0.898_{\text{0.85--0.93}}$ $82.88_{\text{77.1--87.6}}$ $87.00_{\text{82.1--90.3}}$
Common AB $86.19_{\text{81.4--90.0}}$ $84.02_{\text{75.0--89.3}}$ $0.895_{\text{0.84--0.92}}$ $82.91_{\text{77.6--87.8}}$ $86.35_{\text{81.2--89.7}}$
Selected across all except pathology tests (S2) RFECV+LR AB $83.50_{\text{78.5--87.6}}$ $78.64_{\text{68.8--84.8}}$ $0.861_{\text{0.78--0.89}}$ $79.77_{\text{73.8--85.0}}$ $84.21_{\text{79.0--87.9}}$
RFECV+RF MLP $86.25_{\text{81.6--90.1}}$ $77.73_{\text{68.8--83.9}}$ $0.914_{\text{0.84--0.93}}$ $83.14_{\text{77.7--88.5}}$ $89.56_{\text{84.1--91.8}}$
RFECV+GB CB $\textbf{89.23}_{\text{84.9--92.6}}$ $\textbf{83.11}_{\text{74.1--88.4}}$ $0.906_{\text{0.84--0.94}}$ $\textbf{87.12}_{\text{82.0--91.4}}$ $\textbf{90.92}_{\text{86.8--93.7}}$
RFECV+DT DT $87.75_{\text{83.2--91.4}}$ $79.55_{\text{70.5--85.7}}$ $0.913_{\text{0.84--0.94}}$ $85.11_{\text{79.8--90.1}}$ $90.19_{\text{86.2--93.2}}$
RFECV+AB AB $83.50_{\text{78.5--87.6}}$ $78.64_{\text{68.8--84.8}}$ $0.861_{\text{0.78--0.89}}$ $79.77_{\text{73.8--85.0}}$ $84.21_{\text{79.0--87.9}}$
RFECV+ET CB $86.25_{\text{81.6--90.1}}$ $77.73_{\text{68.8--83.9}}$ $0.909_{\text{0.84--0.93}}$ $83.14_{\text{77.7--88.5}}$ $89.56_{\text{84.1--91.8}}$
RFECV+XGB RF $87.17_{\text{82.6--90.8}}$ $79.55_{\text{70.5--85.7}}$ $\textbf{0.921}_{\text{0.86--0.94}}$ $84.24_{\text{79.0--89.3}}$ $89.30_{\text{84.9--92.4}}$
RFECV+CB MLP $86.25_{\text{81.6--90.1}}$ $77.73_{\text{68.8--83.9}}$ $0.914_{\text{0.84--0.93}}$ $83.14_{\text{77.7--88.5}}$ $89.56_{\text{84.1--91.8}}$
LASSO LR $83.50_{\text{78.5--87.6}}$ $78.64_{\text{68.8--84.8}}$ $0.892_{\text{0.83--0.92}}$ $79.77_{\text{73.8--85.0}}$ $84.21_{\text{79.0--87.9}}$
MWUtest CB $84.42_{\text{79.5--88.3}}$ $78.64_{\text{69.6--84.8}}$ $0.896_{\text{0.83--0.93}}$ $80.84_{\text{75.1--86.1}}$ $85.61_{\text{80.2--89.0}}$
Union LGB $84.58_{\text{79.6--88.5}}$ $79.62_{\text{69.6--85.7}}$ $0.898_{\text{0.84--0.93}}$ $80.95_{\text{75.3--86.2}}$ $85.52_{\text{80.3--88.9}}$
Common AB $83.50_{\text{785--87.6}}$ $78.64_{\text{68.8--84.8}}$ $0.861_{\text{0.78--0.89}}$ $79.77_{\text{73.8--85.0}}$ $84.21_{\text{79.0--87.9}}$
@@
Overall, the results in Table 5 highlight that models trained on the selected features from the full feature set (S1) achieve the strongest predictive performance. The highest balanced accuracy of 90.40% (95% CI 86.1%–93.5%) is obtained using a DT model applied to the RFECV with CatBoost selected S1 feature subset (Table 6), achieving the highest sensitivity of 84.85% (95% CI 75.9%–90.2%) and F1 score of 88.77% (95% CI 83.5%–92.6%). Among the individual feature groups, the Medical History (MH) and Socio-Demographic (SD) sets demonstrate competitive predictive ability, with balanced accuracies of 78.87% (95% CI 73.5%–83.2%) and 73.27 (95% CI 67.7%–78.3%)%, respectively. Notably, the combined SD-LH-MH set outperforms all individuals and combined groups, achieving a balanced accuracy of 82.93% (95% CI 77.9%–87.0%), which is higher than when clinical examination features are additionally included (i.e., SD-LH-MH-CE).
Best set Features Description Selection approach Best model